OpenStudy (anonymous):

Richard Laurance purchased a new condominium for $225,000. The bank required a $40,000 down payment. Assume a rate of 11.5% on a 30 year mortgage. What is his monthly payment and total interest cost?

OpenStudy (tkhunny):

This is a Present Value problem. 225000 - 40000 = 185000 185000 = Payment*(Present Value Formula) We can build the formula if you like.

OpenStudy (anonymous):

Ok, yeah that would help me

OpenStudy (tkhunny):

You haven't a magic sheet with that on it? Annual Interest Rate: i = 0.115 Assume Monthly Compounding, not quite staated in the problem. Monthly Interest Rate: j = i/12 = 0.115/12 = 0.00958333... Monthly Discount Factor: v = 1/(1+i) = 0.990507635 30 years ==> 360 Payments You need: \(185000 = Payment(v + v^{2} + ... + v^{360}) = Payment\cdot\dfrac{v-v^{361}}{1-v} = Payment\cdot\dfrac{1-v^{360}}{j}\)

OpenStudy (tkhunny):

That last one is \(Payment\cdot\dfrac{1-v^{360}}{j}\)